# Large Model Empowered Metaverse: State-of-the-Art, Challenges and Opportunities

### v1

@@ Abstract
Index Terms:
Metaverse, large models, cloud-edge-end collaborative rendering, diffusion model.
@@ I Introduction
The Metaverse, an immersive virtual space merging physical and digital realities, was first introduced in the 1990s by Neal Stephenson in his science fiction novel Snow Crash. Recently, advances in virtual reality (VR), augmented reality (AR), artificial intelligence (AI), and blockchain have propelled the Metaverse’s evolution from fiction to reality [1]. Recognized as the next-generation Internet paradigm, the Metaverse envisions a decentralized, immersive, and persistent digital ecosystem where users interact, socialize, create, shop, work, and play through their avatars. Beyond gaming and entertainment, the Metaverse spans diverse applications, including education, healthcare, commerce, and arts, offering new forms of digital living experiences [2]. The Metaverse also aims to establish an open, interconnected digital economy, enabling seamless transfer of assets and identities across various virtual platforms, fostering a diverse, inclusive, and innovation-driven digital society. This transformative digital landscape promises to reshape how people live, interact, and conduct business, ushering in a new era of interconnected, immersive experiences in a blended reality. As reported by P&S Intelligence [^1] https://www.psmarketresearch.com/market-analysis/metaverse-market, the market size was valued at approximately \$129 billion in 2024 and is projected to grow at a compound annual growth rate (CAGR) of 44%, potentially reaching \$1157 billion by 2030.
@@
Large models, particularly large language models (LLMs) and large vision models (LVMs), characterized by their billions of parameters, have witnessed remarkable progress due to recent advancements in generative AI (GAI) algorithms, increased computational capabilities, and availability of Internet-scale datasets [3]. Pioneering large models such as OpenAI’s GPT-4o, Google’s PaLM 2, Microsoft’s Copilot, and OpenAI’s Sora have established a foundation for creating highly sophisticated AI systems capable of understanding and generating human-like text, images, and other media. In the Metaverse, large models play a transformative role, significantly enhancing its content richness, interaction immdersiveness, quality-of-experience (QoE), and quality-of-service (QoS) [2, 4, 5]. Tech giants including Microsoft, Google, OpenAI, Tencent, Meta, NVIDIA, and Alibaba are utilizing large models and GAI to enhance the Metaverse’s capabilities, including virtual environments, digital avatars, interactive VR, 3D digital humans, digital twins, and immersive experiences.
@@
Large models’ powerful data processing and generation abilities contribute to generation of rich, multimodal content, including text, video, images, audio, and 3D interactive objects, which creates immersive and engaging virtual worlds [2]. They enable personalized avatars, interactive NPCs, and digital humans with lifelike behaviors and nuanced emotional expressions, enriching user interactions in virtual environments. Additionally, large models facilitate the construction of intricate virtual worlds by producing accurate digital twins, customized scenes, and context-aware feedback tailored to user interests [4, 5], which enhances engagement and QoE. With their zero/few-shot learning and continuous learning capabilities [6], large models optimize Metaverse performance through intelligent resource allocation and adaptive rendering techniques, improving network efficiency and QoS [7]. Despite these advancements, Metaverse systems powered by large models still face challenges: 1) limited scalability in virtual world construction; 2) constrained responsiveness in rendering; 3) poor adaptability in dynamic environments.
@@ II-A Overview of Large Model Empowered Metaverse
They can also be classified into discriminant and generative models based on their functionality.
  • Discriminant models such as segment anything model (SAM) [^2] https://segment-anything.com/ learn the conditional probability distribution within data and guide decision-making based on data insights, which are widely applied in areas such as intelligent recommendation and risk management.
  • Generative models (e.g., variational autoencoder and diffusion models) learn the joint probability distribution of data and generate new data samples based on learned patterns [8], which are widely used for text generation, image synthesis, code generation, and style transfer.
@@
  • Intent understanding: By combining short-term memory with long-term user knowledge, LLMs can accurately interpret user behavior and intent, supporting adaptive, personalized interactions [8]. This enables avatars to mirror users’ intentions and expressions, making social interactions more lifelike.
@@
2) Perception: Large models enhance the Metaverse’s capacity to understand and interpret real-world and simulated environments.
@@
TABLE I: Academic and Industrial Advances of Large Model-Empowered Metaverse
Ref./Company Benefit Metaverse Application Used Large Model Type of Large Model
Huynh-The et al. [7] Comprehensive AI potential in CV, NLP, digital twins, etc Virtual world construction, gaming, healthcare, smart cities Transformers, Diffusion, GANs, Autoencoders Generative
Chamola et al. [2] Enhances text, image, video, audio, 3D content generation Content creation for Metaverse environments GAI models Generative
@@
Xu et al. [4] Realistic traffic simulations Vehicular mixed-reality Metaverse TSDreamBooth Generative
@@
Google (Project Starline) Realistic 3D communication 3D virtual environments GAI models Generative
@@
NVIDIA (Omniverse) Digital twin creation Simulation, immersive experiences GAI models Generative
@@ II-B State-of-the-Arts
Leading technology companies are harnessing large model technologies to reshape Metaverse applications, fueling innovation and enhancing user engagement. Microsoft integrates AI into its Mesh platform [^3] https://www.uctoday.com/collaboration/microsoft-to-introduce-ai-optimised-avatars-to-mesh-for-teams/, enabling AI-driven avatars for more interactive virtual experiences. Google’s Project Starline [^4] https://starline.google/ leverages GAI to construct realistic 3D environments and immersive communication spaces. OpenAI’s DALL-E [^5] https://openai.com/index/dall-e-3/ powers the creation of unique digital assets, such as artwork and environmental designs, supporting creative customization in virtual worlds. Tencent focuses on AI-driven digital humans and virtual companions [^6] https://www.tencentcloud.com/products/ivh, enriching user interactions in virtual environments. Meta employs large models in Horizon Worlds [^7] https://developers.meta.com/horizon to generate new virtual worlds and personalize experiences, enhancing social VR interactions. NVIDIA’s Omniverse platform [^8] https://www.nvidia.com/en-us/omniverse/ utilizes GAI to create digital twins, facilitating simulations and immersive experiences. Alibaba integrates AI-generated avatars and virtual assistants [^9] https://www.alibabacloud.com/blog/alibaba-clouds-new-ai-tools-spotted-at-apsara-2023_600550 to transform virtual shopping, guiding users in interactive 3D spaces.
@@ III-A Low Scalability in Virtual World Construction
Large models are essential for rendering large-scale, immersive, and highly realistic Metaverse scenes, but their enormous computational demands often result in significant scalability challenges. Current Metaverse platforms, such as Roblox, Fortnite, and Horizon Worlds, while attracting tens of millions of active users daily, struggle to support efficient interactions among large numbers of users. These platforms manage their vast user bases by dividing them into numerous smaller spaces or servers, where each server hosts only a limited number of users. For instance, Horizon Worlds can accommodate a maximum of 20 users interacting simultaneously per space and Fortnite’s sandbox creative mode can only support 16 players per space [^10] https://www.designboom.com/technology/mark-zuckerberg-metaverse-horizon-worlds-avatar-virtual-reality-08-26-2022/, that falls far short of the vision of billions of participants in the future Metaverse.
@@ III-B Constrained Responsiveness in Metaverse Rendering
Current Metaverse rendering techniques often exhaustively render all potentially relevant scenes for each user without considering users’ mobility patterns or specific intentions. This indiscriminate rendering leads to significant computational overhead and resource inefficiency, as many scenes are rendered but remain unused or unnoticed, ultimately degrading overall system performance.
To improve efficiency, integrating user mobility analysis and intention prediction into rendering processes is essential. By employing advanced algorithms that track and analyze user movement patterns, the Metaverse can predict future actions and dynamically adjust rendering priorities to focus on the most relevant areas. For example, if an avatar’s likely path and interactions are anticipated, the system can pre-render only those specific regions, thereby optimizing rendering resources and reducing latency. Additionally, such predictive rendering should be adaptive in real-time to dynamic user behaviors.
@@ III-C Low Adaptability in Dynamic Metaverse Environments
It is essential to develop adaptive scene rendering techniques that can generate personalized and fine-grained content tailored to avatars’ interests, preferences, and context. For instance, avatars who prefer highly interactive elements can be provided with high-resolution, detailed scenes that enhance engagement, while those focused more on social interactions may receive simplified renderings that prioritize avatar clarity and response speed.
@@ IV-A Cloud-Edge-End Collaborative Metaverse Architecture
  • At the top layer, the cloud server serves as the primary storage hub for users’ virtual assets, identities, and other critical data. The cloud’s immense computational power supports large-scale data processing and storage, handling non-time-sensitive tasks and ensuring consistent global access to users’ digital resources.
  • The middle layer consists of edge servers, which are geographically distributed and each manages a sub-Metaverse. Edge servers handle computational tasks that require low latency but need more processing power than the user’s device can provide [10]. By rendering closer to the user, edge servers significantly reduce latency and improve the overall responsiveness of virtual experience.
  • At the bottom layer, users enter the Metaverse through devices such as VR headsets and head-mounted displays and known as “avatars”. Although these devices have limited computational capabilities compared to cloud and edge servers, they are responsible for rendering real-time, low-latency interactions critical for immersion.
@@
  • Foreground interactions refer to user’s immediate actions, such as movements, gestures, and interactions with virtual objects or other avatars. They are less computationally intensive but require minimal latency to ensure smooth, responsive user experience. Typically, foreground interactions are rendered locally on the user’s device, such as a VR headset, which allows for real-time, low-latency feedback that enhances the sense of immersion.
  • Background environment encompasses the broader virtual space where users engage in social activities, exploration, or interactions with dynamic surroundings. They often resembling a virtual panoramic “map” of the virtual world and require a larger rendering area with rich visual details, requiring extensive computational resources.
System Workflow: As depicted in Fig. 2, when a user enters a Metaverse environment, their device handles the rendering of real-time interactions (step 1), such as interacting with other avatars or virtual objects. Simultaneously, large models deployed on the cloud and edge servers cooperatively render background maps (step 2). After that, the high-fidelity background is transmitted to user’s device (step 3), where it synchronizes with the locally rendered foreground interactions (step 4) to create cohesive, immersive experience, based on user’s current field of view (FoV). This process involves several steps: (i) the user requests background rendering, (ii) the cloud and edge collaborate to render and encode the panorama, (iii) the transmission to the user device, and (iv) the device decodes and integrates the background with the foreground interaction.
We utilize a hierarchical game to analyze the optimal strategies of edge nodes and the cloud. The cloud’s utility function is defined as the difference between its revenue from leasing resources and resource costs, while the utility function of edge node is defined as the difference between the resource satisfaction of rendering task and payment. Specifically, tier-1 is a sequential game between the cloud and edge nodes, while tier-2 is a non-cooperative game among edge nodes. Using backward induction, we first analyze the optimal rendering resource demand strategy for each edge node by calculating the best response functions of all other edge nodes, which leads to the Nash equilibrium strategy for each edge node in this non-cooperative game (step 2-1). Then, by substituting optimal strategies of edge nodes into the cloud’s utility function, we derive the cloud’s optimal resource pricing strategy via an iterative algorithm based on gradient descent due to the difficulty in obtaining its explicit expression (step 2-2).
@@ IV-B Mobility-Aware Metaverse Pre-Rendering Mechanism
Rendering the background environment in the Metaverse involves the above four steps, which can be time-consuming and decrease frame rates. Additionally, misalignment between rendered scenes and users’ movements can negatively impact overall QoE. To tackle these challenges, we employ a pre-rendering technique that utilizes panorama compression while leveraging user mobility and spatial similarity.
@@ IV-C Diffusion Model Based Personalized Metaverse Rendering Strategy
Phase 1: Preference-oriented forward diffusion. As depicted in Fig. 3(a), we first model user’s multi-faceted behaviors in the Metaverse, with the historical behavior sequence representing his/her Metaverse preferences. Specifically, we decode data packets from the user’s previous Metaverse activities, assembling interaction data (e.g., visited areas and interacted items), latency data (e.g., network and buffering latency), and fluency data (e.g., refresh rate and frame rate) into a multi-variable sequence M, representing the Meteverse user’s preferences. Next, we apply a forward diffusion process to incrementally perturb M into M’ via controlled transformations, adding stochastic noise to simulate potential shifts in user interests and context. Each iteration introduces randomness, enabling the diffusion model to explore a broad range of potential preferences.
Phase 2: Resource-aware reverse denoising. As shown in Fig. 3(b), the reverse denosing phase is designed to progressively reconstruct user’s original preferences M from the distorted preferences M’ generated during the forward diffusion phase, learning to forecast user’s preferences when rendering new scenes in the Meteverse. As user’s preferences in scene quality is often constrained by device status (e.g., CPU/GPU load, battery level) and network performance(e.g., bandwidth fluctuation), we set up resource constraints S by monitoring user’s device status and network performance in real-time. For each reverse denosing iteration, Attention U-Net [13] is adopted to learn the conditional probabilities under resource constraints S, where S serves as the additional input to guide the reverse denoising of M’, allowing the diffusion model to adaptively learn user’s preferences based on resource constraints.
Phase 3: Personalized real-time rendering. As illustrated in Fig. 3(c), the diffusion model first utilizes user’s historical data to conduct offline training for preferences learning, then leverages Metaverse real-time data to perform online rendering tailed to user’s preferences and resource constraints. During offline training, we aim to predict noise of user’s distorted preferences M’ in the reverse denoising phase and minimizes the mean square error between Gaussian noise and the predicted ones. After offline training, the diffusion model can conduct online rendering. For a given new scene, we perform forward diffusion and reverse denoising to reconstruct the interaction probabilities of non-interacted items in the new scene. By utilizing these predicted interaction probabilities, we dynamically adjust the rendering focus in real time, ensuring high-quality visuals in areas preferred by the user while adapting scene details to align with personalized interests and resource constraints.

### v2

@@ I Introduction
The Metaverse, an immersive virtual space merging physical and digital realities, was first introduced in the 1990s by Neal Stephenson in his science fiction novel Snow Crash. Recently, advances in virtual reality (VR), augmented reality (AR), artificial intelligence (AI), and blockchain have propelled the Metaverse's evolution from fiction to reality [1]. Recognized as the next-generation Internet paradigm, the Metaverse envisions a decentralized, immersive, and persistent digital ecosystem where users interact, socialize, create, shop, work, and play through their avatars. Beyond gaming and entertainment, the Metaverse spans diverse applications, including education, healthcare, commerce, and arts, offering new forms of digital living experiences [2]. The Metaverse also aims to establish an open, interconnected digital economy, enabling seamless transfer of assets and identities across various virtual platforms, fostering a diverse, inclusive, and innovation-driven digital society. This transformative digital landscape promises to reshape how people live, interact, and conduct business, ushering in a new era of interconnected, immersive experiences in a blended reality. As reported by P&S Intelligence [^1] https://www.psmarketresearch.com/market-analysis/metaverse-market, the market size was valued at approximately \$129 billion in 2024 and is projected to grow at a compound annual growth rate (CAGR) of 44%, potentially reaching \$1157 billion by 2030.
@@
Large models, particularly large language models (LLMs) and large vision models (LVMs), characterized by their billions of parameters, have witnessed remarkable progress due to recent advancements in generative AI (GAI) algorithms, increased computational capabilities, and availability of Internet-scale datasets [3]. Pioneering large models such as OpenAI's GPT-4o, Google's PaLM 2, Microsoft's Copilot, and OpenAI's Sora have established a foundation for creating highly sophisticated AI systems capable of understanding and generating human-like text, images, and other media. In the Metaverse, large models play a transformative role, significantly enhancing its content richness, interaction immdersiveness, quality-of-experience (QoE), and quality-of-service (QoS) [2, 4, 5]. Tech giants including Microsoft, Google, OpenAI, Tencent, Meta, NVIDIA, and Alibaba are utilizing large models and GAI to enhance the Metaverse's capabilities, including virtual environments, digital avatars, interactive VR, 3D digital humans, digital twins, and immersive experiences.
@@
Large models' powerful data processing and generation abilities contribute to generation of rich, multimodal content, including text, video, images, audio, and 3D interactive objects, which creates immersive and engaging virtual worlds [2]. They enable personalized avatars, interactive NPCs, and digital humans with lifelike behaviors and nuanced emotional expressions, enriching user interactions in virtual environments. Additionally, large models facilitate the construction of intricate virtual worlds by producing accurate digital twins, customized scenes, and context-aware feedback tailored to user interests [4, 5], which enhances engagement and QoE. With their zero/few-shot learning and continuous learning capabilities [6], large models optimize Metaverse performance through intelligent resource allocation and adaptive rendering techniques, improving network efficiency and QoS [7]. Despite these advancements, Metaverse systems powered by large models still face challenges: 1) limited scalability in virtual world construction (i.e., how to support large-scale, immersive multi-user interactions under high computational cost of large model-driven scene rendering); 2) constrained responsiveness in rendering (i.e., how to improve rendering responsiveness by integrating user mobility and intention prediction into real-time scene adaptation); 3) poor adaptability in dynamic environments (i.e., how to enable personalized, fine-grained scene rendering that adapts to diverse user preferences, device capabilities, and dynamic interaction contexts).
@@ II-A Overview of Large Model Empowered Metaverse
  • Intent understanding: By combining short-term memory with long-term user knowledge, LLMs can accurately interpret user behavior and intent, supporting adaptive, personalized interactions [8]. This enables avatars to mirror users' intentions and expressions, making social interactions more lifelike.
@@
2) Perception: Large models enhance the Metaverse's capacity to understand and interpret real-world and simulated environments.
@@
TABLE I: Academic and Industrial Advances of GAI and Large Model Empowered Metaverse
Ref./Company Benefit Metaverse Application Used GAI Model Type of Large Model
Huynh-The et al. [7] Comprehensive AI potential in CV, NLP, digital twins, etc Virtual world construction, gaming, healthcare, smart cities Transformers, Diffusion, GANs, Autoencoders
Chamola et al. [2] Enhances text, image, video, audio, 3D content generation Content creation for Metaverse environments GAI models
@@
Xu et al. [4] Realistic traffic simulations Vehicular mixed-reality Metaverse TSDreamBooth
@@
Google (Project Starline) Realistic 3D communication 3D virtual environments GAI models
@@
NVIDIA (Omniverse) Digital twin creation Simulation, immersive experiences GAI models
@@ II-B State-of-the-Arts
Leading technology companies are harnessing large model technologies to reshape Metaverse applications, fueling innovation and enhancing user engagement. Microsoft integrates AI into its Mesh platform [^3] https://www.uctoday.com/collaboration/microsoft-to-introduce-ai-optimised-avatars-to-mesh-for-teams/, enabling AI-driven avatars for more interactive virtual experiences. Google's Project Starline [^4] https://starline.google/ leverages GAI to construct realistic 3D environments and immersive communication spaces. OpenAI's DALL-E [^5] https://openai.com/index/dall-e-3/ powers the creation of unique digital assets, such as artwork and environmental designs, supporting creative customization in virtual worlds. Tencent focuses on AI-driven digital humans and virtual companions [^6] https://www.tencentcloud.com/products/ivh, enriching user interactions in virtual environments. Meta employs large models in Horizon Worlds [^7] https://developers.meta.com/horizon to generate new virtual worlds and personalize experiences, enhancing social VR interactions. NVIDIA's Omniverse platform [^8] https://www.nvidia.com/en-us/omniverse/ utilizes GAI to create digital twins, facilitating simulations and immersive experiences. Alibaba integrates AI-generated avatars and virtual assistants [^9] https://www.alibabacloud.com/blog/alibaba-clouds-new-ai-tools-spotted-at-apsara-2023_600550 to transform virtual shopping, guiding users in interactive 3D spaces.
@@ III-A Low Scalability in Virtual World Construction
Large models are essential for rendering large-scale, immersive, and highly realistic Metaverse scenes, but their enormous computational demands often result in significant scalability challenges. Current Metaverse platforms, such as Roblox, Fortnite, and Horizon Worlds, while attracting tens of millions of active users daily, struggle to support efficient interactions among large numbers of users. These platforms manage their vast user bases by dividing them into numerous smaller spaces or servers, where each server hosts only a limited number of users. For instance, Horizon Worlds can accommodate a maximum of 20 users interacting simultaneously per space and Fortnite's sandbox creative mode can only support 16 players per space [^10] https://www.designboom.com/technology/mark-zuckerberg-metaverse-horizon-worlds-avatar-virtual-reality-08-26-2022/, that falls far short of the vision of billions of participants in the future Metaverse.
@@ III-B Constrained Responsiveness in Metaverse Rendering
Current Metaverse rendering techniques often exhaustively render all potentially relevant scenes for each user without considering users' mobility patterns or specific intentions. This indiscriminate rendering leads to significant computational overhead and resource inefficiency, as many scenes are rendered but remain unused or unnoticed, ultimately degrading overall system performance.
To improve efficiency, integrating user mobility analysis and intention prediction into rendering processes is essential. By employing advanced algorithms that track and analyze user movement patterns, the Metaverse can predict future actions and dynamically adjust rendering priorities to focus on the most relevant areas. For example, if an avatar's likely path and interactions are anticipated, the system can pre-render only those specific regions, thereby optimizing rendering resources and reducing latency. Additionally, such predictive rendering should be adaptive in real-time to dynamic user behaviors.
@@ III-C Low Adaptability in Dynamic Metaverse Environments
It is essential to develop adaptive scene rendering techniques that can generate personalized and fine-grained content tailored to avatars' interests, preferences, and context. For instance, avatars who prefer highly interactive elements can be provided with high-resolution, detailed scenes that enhance engagement, while those focused more on social interactions may receive simplified renderings that prioritize avatar clarity and response speed.
@@ IV-A Cloud-Edge-End Collaborative Metaverse Architecture
  • At the top layer, the cloud server serves as the primary storage hub for users' virtual assets, identities, and other critical data. The cloud's immense computational power supports large-scale data processing and storage, handling non-time-sensitive tasks and ensuring consistent global access to users' digital resources.
  • The middle layer consists of edge servers, which are geographically distributed and each manages a sub-Metaverse. Edge servers handle computational tasks that require low latency but need more processing power than the user's device can provide [10]. By rendering closer to the user, edge servers significantly reduce latency and improve the overall responsiveness of virtual experience.
  • At the bottom layer, users enter the Metaverse through devices such as VR headsets and head-mounted displays and known as ``avatars''. Although these devices have limited computational capabilities compared to cloud and edge servers, they are responsible for rendering real-time, low-latency interactions critical for immersion.
@@
  • Foreground interactions refer to user's immediate actions, such as movements, gestures, and interactions with virtual objects or other avatars. They are less computationally intensive but require minimal latency to ensure smooth, responsive user experience. Typically, foreground interactions are rendered locally on the user's device, such as a VR headset, which allows for real-time, low-latency feedback that enhances the sense of immersion.
  • Background environment encompasses the broader virtual space where users engage in social activities, exploration, or interactions with dynamic surroundings. They often resembling a virtual panoramic ``map'' of the virtual world and require a larger rendering area with rich visual details, requiring extensive computational resources.
System Workflow: As depicted in Fig. 2, when a user enters a Metaverse environment, their device handles the rendering of real-time interactions (step 1), such as interacting with other avatars or virtual objects. Simultaneously, large models deployed on the cloud and edge servers cooperatively render background maps (step 2). After that, the high-fidelity background is transmitted to user's device (step 3), where it synchronizes with the locally rendered foreground interactions (step 4) to create cohesive, immersive experience, based on user's current field of view (FoV). This process involves several steps: (i) the user requests background rendering, (ii) the cloud and edge collaborate to render and encode the panorama, (iii) the transmission to the user device, and (iv) the device decodes and integrates the background with the foreground interaction.
We utilize a hierarchical game to analyze the optimal strategies of edge nodes and the cloud. The cloud's utility function is defined as the difference between its revenue from leasing resources and resource costs, while the utility function of edge node is defined as the difference between the resource satisfaction of rendering task and payment. Specifically, tier-1 is a sequential game between the cloud and edge nodes, while tier-2 is a non-cooperative game among edge nodes. Using backward induction, we first analyze the optimal rendering resource demand strategy for each edge node by calculating the best response functions of all other edge nodes, which leads to the Nash equilibrium strategy for each edge node in this non-cooperative game (step 2-1). Then, by substituting optimal strategies of edge nodes into the cloud's utility function, we derive the cloud's optimal resource pricing strategy via an iterative algorithm based on gradient descent due to the difficulty in obtaining its explicit expression (step 2-2).
@@ IV-B Mobility-Aware Metaverse Pre-Rendering Mechanism
Rendering the background environment in the Metaverse involves the above four steps, which can be time-consuming and decrease frame rates. Additionally, misalignment between rendered scenes and users' movements can negatively impact overall QoE. To tackle these challenges, we employ a pre-rendering technique that utilizes panorama compression while leveraging user mobility and spatial similarity.
@@ IV-C Diffusion Model Based Personalized Metaverse Rendering Strategy
Phase 1: Preference-oriented forward diffusion. As depicted in Fig. 3(a), we first model user's multi-faceted behaviors in the Metaverse, with the historical behavior sequence representing his/her Metaverse preferences. Specifically, we decode data packets from the user's previous Metaverse activities, assembling interaction data (e.g., visited areas and interacted items), latency data (e.g., network and buffering latency), and fluency data (e.g., refresh rate and frame rate) into a multi-variable sequence M, representing the Meteverse user's preferences. Next, we apply a forward diffusion process to incrementally perturb M into M' via controlled transformations, adding stochastic noise to simulate potential shifts in user interests and context. Each iteration introduces randomness, enabling the diffusion model to explore a broad range of potential preferences.
Phase 2: Resource-aware reverse denoising. As shown in Fig. 3(b), the reverse denosing phase is designed to progressively reconstruct user's original preferences M from the distorted preferences M' generated during the forward diffusion phase, learning to forecast user's preferences when rendering new scenes in the Meteverse. As user's preferences in scene quality is often constrained by device status (e.g., CPU/GPU load, battery level) and network performance(e.g., bandwidth fluctuation), we set up resource constraints S by monitoring user's device status and network performance in real-time. For each reverse denosing iteration, Attention U-Net [13] is adopted to learn the conditional probabilities under resource constraints S, where S serves as the additional input to guide the reverse denoising of M', allowing the diffusion model to adaptively learn user's preferences based on resource constraints.
Phase 3: Personalized real-time rendering. As illustrated in Fig. 3(c), our proposed diffusion scheme adopts a hybrid offline-online rendering strategy. It first utilizes user's historical data to conduct offline training for preferences learning, then leverages Metaverse real-time data to perform online rendering tailed to user's preferences and resource constraints. During offline training, we aim to predict noise of user's distorted preferences M' in the reverse denoising phase and minimizes the mean square error between Gaussian noise and the predicted ones with step-by-step error analysis, ensuring effectiveness and preference alignment.
After offline training, the diffusion model can conduct online rendering. For a given new scene, we perform forward diffusion followed by skip-step reverse denoising to reconstruct the interaction probabilities of non-interacted items in the new scene, thereby expediting the inference process of the trained diffusion model for real-time personalized rendering. By utilizing these predicted interaction probabilities, we dynamically adjust the rendering focus in real time, ensuring high-quality visuals in areas preferred by the user while adapting scene details to align with personalized interests and resource constraints.